### v1

# industrial control system STIX Structured Threat Information Expression

ICS
industrial control system
STIX
Structured Threat Information Expression
AIS
Automated Indicator Sharing
DOE
Department of Energy
CRISP
Cybersecurity Risk Information Sharing Program
CyOTE
Cybersecurity for the Operational Technology Environment
KEV
Known Exploitable Vulnerability
IOC
indicators of compromise
LLM
large language model
SCO
STIX Cyber-observable Object
CTI
Cyber Threat Intelligence
CPS
Cyber-Physical System
OT
Operational Technology
IT
Information Technology
NIS
Network and Information Systems
PLC
Programmable Logic Controller
CERT
Cyber Emergency Response Teams
TTP
Tactics, Techniques and Procedures
API
Application Programming Interface
OS
Operating System
C2
Command and Control
CRC
Cyclic Redundancy Check
DLL
dynamic-link library
IEC
International Electrotechnical Commission
SCADA
Supervisory Control and Data Acquisition
IOA
Information Object Addresses
PCAP
Packet Capture
CVE
Common Vulnerabilities and Exposures
CWE
Common Weakness Enumeration
ISAC
Information Sharing and Analysis Center
CISA
Cybersecurity and Infrastructure Security Agency
An Evidence-Driven Analysis of Threat Information Sharing Challenges for Industrial Control Systems and Future Directions
Adam Hahn The MITRE Corporation, McLean, VA, USA Rubin Krief Department of Software and Information Systems Engineering, Ben-Gurion University of the Negev, Beer-Sheva, Israel Cyber@BGU, Ben-Gurion University of the Negev, Beer-Sheva, Israel Daniel Rebori-Carretero The MITRE Corporation, McLean, VA, USA Rami Puzis Department of Software and Information Systems Engineering, Ben-Gurion University of the Negev, Beer-Sheva, Israel Cyber@BGU, Ben-Gurion University of the Negev, Beer-Sheva, Israel Aviad Elyashar Cyber@BGU, Ben-Gurion University of the Negev, Beer-Sheva, Israel Department of Computer Science, Shamoon College of Engineering, Beer-Sheva, Israel Nik Urlaub National Renewable Energy Laboratory (NREL)
Abstract
The increasing cyber threats to critical infrastructure highlight the importance of private companies and government agencies in detecting and sharing information about threat activities. Although the need for improved threat information sharing is widely recognized, various technical and organizational challenges persist, hindering effective collaboration. In this study, we review the challenges that disturb the sharing of usable threat information to critical infrastructure operators within the industrial control system (ICS) domain. We analyze three major incidents: Stuxnet, Industroyer, and Triton. In addition, we perform a systematic analysis of 196 procedure examples across 79 MITRE ATT&CK® techniques from 22 ICS-related malware families, utilizing automated natural language processing techniques to systematically extract and categorize threat observables. Additionally, we investigated nine recent ICS vulnerability advisories from the Cybersecurity and Infrastructure Security Agency (CISA) Known Exploitable Vulnerability catalog. Our analysis identified four important limitations in the ICS threat information sharing ecosystem: (i) the lack of coherent representation of artifacts related to ICS adversarial techniques in information sharing language standards (e.g., Structured Threat Information Expression (STIX)); (ii) the dependence on undocumented proprietary technologies; (iii) limited technical details provided in vulnerability and threat incident reports; and (iv) the accessibility of technical details for observed adversarial techniques. This study aims to guide the development of future information-sharing standards, including the enhancement of the cyber-observable objects schema in STIX, to ensure accurate representation of artifacts specific to ICS environments.
Keywords— Threat information sharing cyber threat intelligence reports ATT&CK techniques STIX
1 Introduction
As critical infrastructure increasingly becomes the target of cyber threats, organizations and governments must be able to detect and share information about threat actor activity. Numerous government reports, such as Executive Order 13691 [exec_order_13691] and the Cyberspace Solarium Commission report [cyberspace_solarium_2023], have highlighted the need for enhanced information sharing to safeguard critical infrastructure. Additionally, numerous government programs have been established to facilitate information sharing between the public and private sectors.
Effective information sharing enhances collective situational awareness, enabling quicker responses to emerging threats, and reducing the risk of widespread disruption. However, realizing these benefits requires not only the collection and dissemination of threat information but also ensuring that the shared intelligence contains sufficient technical detail and can be represented in standardized formats that enable operational use by asset owners.
While the need for improved threat information sharing is well-defined, several technical and organizational challenges persist that hinder effective sharing. Prior work has assessed indicators of compromises in ICS, but has not reviewed ATT&CK techniques to determine what observables are required to capture attack behavior.
In this paper, we examine the challenges that hinder the sharing of actionable threat information with asset owners within the ICS domain. This study leverages the MITRE ATT&CK for ICS [mitre_attack_ics] knowledge base, which aggregates cyber threat intelligence reports across known ICS incidents.
Our paper highlights the key challenges by reviewing three well-known incidents: Stuxnet [stuxnet_facts_report], Industroyer [ukraine_power_grid_analysis], and Triton [saudi_arabia_cyberattack]. This analysis aims to identify the adversarial techniques employed by threat actors and to determine the information that asset owners need to implement effective detection mechanisms against these techniques. We also explore nine recent ICS vulnerability advisories with evidence of exploitation as documented in the CISA Known Known Exploitable Vulnerability (KEV) Catalog  [cisa_kev_catalog].
To provide a comprehensive assessment beyond these three high-profile incidents, this study extends the analysis to a systematic examination of 196 procedure examples across 79 techniques from 22 malware families documented in the MITRE ATT&CK for ICS framework. Using a large language model (LLM) to extract and categorize threat observables, we analyzed 361 total observables to understand the breadth of representation and actionability challenges across the entire knowledge base. This comprehensive analysis revealed that only 101 observables had full STIX observable support, 191 had partial support, and 69 lacked any STIX representation. Furthermore, only 87 observables contained actionable technical details sufficient for detection development, while 274 lacked the necessary specificity for operational use.
The primary research questions driving this study are:
  1. 1.
    What are the technical artifacts necessary to detect known ICS adversarial techniques? Can current information-sharing standards adequately represent them? (see Section 4.1).
  2. 2.
    To what extent do existing threat reports provide actionable technical details for detection development? (see Section 4.2.2).
  3. 3.
    To what extent do existing vulnerability advisories provide actionable technical details for detection development? (see Section 5).
  4. 4.
    What are the main challenges and systemic barriers preventing effective threat information sharing in ICS environments? (see Section 6).
The outcome of this analysis identifies four important limitations in the ICS threat information sharing ecosystem that impact the ability of asset owners to effectively use such information. Specifically, this effective information sharing is limited by (i) the lack of coherent representation of artifacts associated with known ICS adversarial techniques in information sharing language standards (e.g., STIX), (ii) the reliance on undocumented and proprietary technologies, especially network protocols without open-source parsers, (iii) the insufficient technical details provided in vulnerability and threat incident reports, and (iv) the availability of technical details across all observed adversarial techniques.
This paper intends to inform the development of future information-sharing standards in ICS, including the expansion of the STIX Cyber-observable Object (SCO) schema. Our study reviews systematically ATT&CK techniques to determine the necessary indicators or observables that represent adversarial behavior across the full spectrum of documented ICS incidents. This approach provides both depth through detailed case study analysis and breadth through a comprehensive systematic assessment of the entire ATT&CK for ICS knowledge base.
2 Related Work
Recent advancements in Cyber Threat Intelligence (CTI) have underscored the complexities of securing ICS and broader Cyber-Physical Systems, particularly in the context of critical infrastructure. ICSs and CPSs often rely on proprietary and specialized protocols that are difficult to monitor and integrate with conventional cybersecurity tools. Krasznay et al. [krasznay2021possibilities] highlighted the critical vulnerabilities in these systems, which stem from their original design priorities focused on safety rather than cybersecurity. The increasing connectivity of ICS and Operational Technology (OT) systems has exposed these vulnerabilities to sophisticated adversaries, including state-sponsored actors, necessitating a paradigm shift in cybersecurity practices. The SeConSys initiative in Hungary represents a significant milestone in addressing these challenges. This initiative facilitates sector-specific threat intelligence sharing and detection through the development of tailored CTI feeds, honeypots that simulate energy-sector environments, and the implementation of the Network and Information Systems (NIS) Directive. These efforts have been instrumental in promoting localized, collaborative approaches to cybersecurity and advancing frameworks for real-time threat information sharing among key stakeholders in the energy sector.
López-Morales et al. [lopez2024securing] extended these contributions by developing innovative methodologies that advance the state-of-the-art in CTI for CPSs. Their work introduces the ICS2 Matrix, an enhanced version of the MITRE ATT&CK framework tailored specifically to ICSs. The ICS2 Matrix incorporates novel attack techniques and mitigation strategies derived from an extensive analysis of Programmable Logic Controller (PLC)-related security research spanning 17 years. Additionally, the development of a high-interaction satellite honeypot provides a groundbreaking tool for simulating realistic adversary interactions, collecting data on emerging threats, and advancing the understanding of satellite-specific attack vectors. This work further proposes the creation of a sandbox environment for connected autonomous vehicles, enabling the simulation and analysis of cyberattacks on autonomous systems to improve resilience. Such contributions underscore the necessity of domain-specific CTI methodologies that address the unique operational and security challenges inherent in different CPSs.
2.1 Empirical Measurements and Quality Assessment of CTI Sharing
At the ecosystem level, empirical studies have begun to quantify the effectiveness and challenges of CTI sharing at scale. Jin et al. [jin2024sharing] provided one of the first comprehensive empirical measurements of CTI sharing effectiveness, analyzing STIX volume, timeliness, coverage, and quality across public sources. Their findings reveal significant variations in feed quality and highlight the challenges that organizations encounter while operating in a shared intelligence environment. This empirical foundation complements broader systematic reviews that examine how organizations actually use CTI for security decision-making [ainslie2023cyber] and identify persistent barriers to effective CTI sharing, including governance issues, misaligned incentives, and insufficient standardization [alaeifar2024current].
Quality and trust of shared feeds remain persistent issues across the CTI ecosystem. Griffioen et al. [griffioen2019quality] conducted systematic quality evaluations of cyber threat intelligence feeds, revealing inconsistent timeliness and specificity that can undermine the operational value of shared intelligence. Similarly, Schaberreiter et al. [schaberreiter2019quantitative] developed quantitative frameworks for evaluating trust in CTI quality, demonstrating that trust and quality metrics vary significantly across different sources and feed types. Together, these studies highlight the need for more rigorous quality assurance mechanisms in CTI sharing initiatives.
2.2 Standardization and Technical Implementation
Standardization efforts play a crucial role in making threat observables portable and actionable across different organizational contexts. STIX 2.1, approved as an OASIS Standard (on 10 June 2021), serves as the exchange format used across many CTI producers and consumers [stix_v21_official]. However, challenges remain in effectively representing complex ICS-specific artifacts within existing STIX objects. Zych and Mavroeidis [zych2022enhancing] have proposed enhancements to MITRE ATT&CK representations by STIX, to improve filtering and prioritization tasks, while RFC 9424 [rfc9424] provides important clarification on the capabilities and limitations of IOCs, reinforcing the value of higher-fidelity observables for creating durable detection capabilities.
Recent advances in natural language processing have introduced new approaches for extracting threat intelligence from unstructured reports. While some pipelines pursue full STIX coverage through classical NLP techniques combined with external knowledge bases [marchiori2023stixnet], alternative approaches rely on large language models to extract observables directly from text without requiring extensive knowledge graphs. These automated extraction techniques show promise for scaling CTI processing, though systematic evaluation reveals significant limitations and failure modes. Alam et al. [alam2024ctibench] developed CTIBench, a comprehensive benchmark specifically designed to evaluate LLM performance in cyber threat intelligence tasks across multiple cognitive dimensions, including memorization, understanding, problem-solving, and reasoning. Their evaluation of five state-of-the-art models revealed concerning patterns: all models struggled significantly with tasks requiring domain expertise, with even the best-performing model (GPT-4) achieving only 71% accuracy on foundational CTI knowledge questions. More critically, their analysis identified systematic failure modes, including overestimation of threat severity scores, difficulty with out-of-knowledge cutoff information, and particularly poor performance on questions related to mitigation strategies and adversarial tools. These findings highlight substantial gaps in LLM reliability for cybersecurity applications.
Fieblinger et al. [fieblinger2024actionable] developed a comprehensive methodology for automating the extraction of actionable CTI using LLMs for generating Knowledge Graphs. Their approach systematically evaluates multiple open-source LLMs (including Llama 2 series, Mistral 7B Instruct, and Zephyr) across different extraction techniques, such as few-shot prompt engineering, guidance frameworks, and fine-tuning to optimize the extraction of meaningful triples from CTI texts. Notably, their methodology demonstrates that guidance frameworks significantly improve extraction performance beyond what is achievable with prompt engineering alone, while fine-tuning proves essential for extracting triples that adhere to specified ontologies. The extracted triples are subsequently used to construct knowledge graphs that provide structured, queryable representations of threat intelligence, facilitating downstream applications such as link prediction and threat analysis.
The development of domain-specific evaluation datasets has become increasingly important for assessing automated threat intelligence processing. Bhusal et al. [bhusal2024secure] developed the SECURE benchmark using MITRE ATT&CK as a foundation for generating cybersecurity LLM performance evaluation across extraction, understanding, and reasoning tasks. Their approach demonstrates the potential of leveraging structured threat intelligence frameworks like ATT&CK for creating systematic evaluation datasets. However, the reliability of LLM-based approaches in cybersecurity contexts remains a critical concern. Their comprehensive evaluation of seven state-of-the-art models revealed significant limitations, including hallucinations when models lack current information, difficulty with out-of-distribution detection, and substantial performance variations across different cybersecurity task types. Importantly, their findings demonstrate that while LLMs show promise for cybersecurity advisory roles, closed-source models consistently outperform open-source alternatives, particularly in problem-solving and out-of-distribution scenarios. These results underscore the importance of meticulous validation and human oversight when deploying LLMs for threat intelligence processing, particularly in critical infrastructure environments where inaccurate information can have severe consequences.
2.2.1 ATT&CK Framework Application in OT Environments
For technique-level grounding in OT environments, the MITRE ATT&CK framework and its ICS extension provide structured approaches to understanding adversary behaviors [mitre_attack_design, mitre_attack_ics_design]. Afenu et al. [afenu2024industrial] have leveraged ATT&CK to validate ICS defense mechanisms and testbed implementations. Choi et al. [choi2021probabilistic] generated realistic OT attack sequences based on analysis of real-world datasets. These applications demonstrate the framework’s utility in bridging the gap between theoretical threat models and practical security implementations.
The existing literature highlights several recurring challenges in CTI, including limited interoperability of detection tools, insufficient standardization of threat-sharing protocols, and a lack of actionable intelligence for specific adversarial behaviors. While Krasznay et al. [krasznay2021possibilities] emphasized the importance of legislative and collaborative frameworks, López-Morales et al. [lopez2024securing] focused on advancing technical methodologies to enhance the granularity and applicability of CTI. Both approaches converge on the need for tailored threat intelligence that integrates sector-specific requirements into practical and effective security solutions. Building on these foundational works, this study aims to address critical gaps in the ICS threat information-sharing ecosystem. Our research specifically focuses on expanding the cyber observable objects schema within the STIX standard to ensure the accurate representation of ICS-specific artifacts and adversarial techniques. Unlike prior studies that primarily concentrated on domain-specific or geographically localized advancements, this work seeks to generalize CTI enhancements across diverse ICS environments. By integrating detailed technical insights into standardized threat-sharing protocols, this study aspires to bridge the gap between theoretical threat models and actionable intelligence, thereby facilitating more effective collaboration and situational awareness across ICS stakeholders.
3 Key Roles and Functions in Sharing Threat/Vulnerability Information
This section provides a brief overview of the ICS information sharing ecosystem, including key organizations, information flows, artifacts, and systems that support the sharing of threat and vulnerability information. Specifically, we focus on the needs of the asset owners who rely on this information to support security operations and detection efforts (see Figure 1).
Figure 1: Overview of Threat/Vulnerability Information Sharing Roles and Technical Capabilities. Solid boxes represent stakeholders. Dashed boxes represent components. Solid arrows represent common existing information flows.
Typically, governments facilitate information sharing by providing reports with threat or vulnerability information to critical infrastructures and other organizations [gao_cybersecurity_2023]. The National Cyber Emergency Response Teams (CERT) is an important information-sharing hub, as it has unique access to threat information through the facilitation of information-sharing programs with both federal, defense, and intelligence agencies, as well as partnerships with private sector organizations.
Security vendors provide tools/platforms and information to support the detection and mitigation of cyber threats [splunk_tdir, crowdstrike_tdir]. Many of these vendors search for vulnerabilities and threat actor activity to develop analytics or rules to be incorporated into detection tools. Some vendors provide machine-readable threat information feeds, which can be directly consumed by security products, including those used to monitor for cyber threats. Further, many vendors obtain data from their detection infrastructures deployed at asset owner sites or collect information while supporting incident response efforts.
ICS vendors typically support information sharing efforts through the disclosure of security advisories and patches when a vulnerability is discovered in their software or devices [ics_patch_tuesday]. Most importantly, ICS vendors have unique knowledge regarding the technical details for their products, which are typically based on proprietary technology.
Asset owners are responsible for managing the risk presented by the cyber threats. The larger organizations that maintain critical infrastructures typically have dedicated security operations teams, the smaller commercial asset owners commonly rely on their national CERT [Cybersecurity_Guide_for_SMEs]. They may have a person partially employed in security operations to comply with regulations. The security team of an asset owner is responsible for monitoring their systems for threats and reporting to their auditors or the CERT.
Security operations teams depend on several different security products to support this, including detection platforms that monitor ICS devices and networks for threat activities based on various detection rules and analytics. The security operations teams depend on governments, ICS vendors, and security vendors to provide information about a range of vulnerabilities and threats.
In addition to the previously identified organizations, other entities contribute to information-sharing capabilities. For example, various organizations have provided open-source tools to parse network traffic, extract security-related data, and apply analytics to detect adversary behaviors shared with the organizations. Examples of key open-source efforts supporting the dissemination and detection of threat behaviors are included below.
  • Standardized threat information languages: One key contribution to make threat information sharing more effective is the development of standard languages. For example, STIX standard provides a JSON format to communicate threat behaviors. This includes definitions of SCOs [stix_v21_official] for specific types of artifacts that can be used to detect malicious activities or collect information about them. Other formats include OpenIOC [^1] https://www.scribd.com/document/653654203/An-Introduction-to-OpenIOC and MISP XML [^2] https://www.misp-project.org/.
  • Detection Rules/Analytics: Various security vendors provide rules or analytics that can be integrated within a detection platform to enable the automated detection of specific behaviors. While many commercial detection platforms consume their own proprietary feed of detection rules or analytics, others rely on open-source formats. Examples include the NSA ELITEWOLF [^3] https://github.com/nsacyber/ELITEWOLF and CISA ATT&CK-based Control-system Indicator Detection (ACID), [^4] https://github.com/cisagov/ACID and various repositories of YARA rules.
  • ICS Protocol Parsers: Many detection platforms lack inherent support for ICS protocols; therefore, dedicated parsers for ICS protocols have been developed to enable the monitoring of ICS networks. Programs like CISA’s Industrial Control Systems Network Protocol Parsers [^5] https://github.com/cisagov/ICSNPP have defined several parsers for key ICS protocols for Zeek, while Cisco Talos has also now identified Snort inspectors for specific ICS protocols. [^6] https://blog.talosintelligence.com/ics-protocol-coverage-snort-3/
Requirements for Threat Detection: To ensure an asset owner can detect a specific threat behavior, they must either (i) obtain a detection rule or analytic to support that detection, or (ii) develop a detection based on known technical data about that threat and the associated technology it targets. Accordingly, the following two requirements are defined to support the analysis of information sharing effectiveness and identify associated gaps.
  • Requirement #1 - The information should include a detection rule/analytics or structured cyber observables that could be directly integrated within a detection platform.
  • Requirement #2 - The information should include sufficient technical details to enable the development of a detection rule/analytic. However, to ensure the end-user has sufficient information to establish a detection, the following additional sub-requirements must be met:
    1. a)
      Technical details about the threat activity, including the adversarial technique used, and the specific artifacts manipulated during the threat activity.
    2. b)
      A thorough understanding of the underlying technology and protocols targeted by the threat is required to develop an effective detection. This is particularly relevant when a threat targets a vendor’s proprietary technology, especially if the technology is not adequately documented. Developing detectors for undocumented third-party technologies is a laborious and unreliable task.
4 Threat Information Sharing Analysis
This section reviews the procedure examples defined in MITRE ATT&CK for ICS associated with the Stuxnet, Industroyer, and Triton events. While the ATT&CK for ICS does not represent any single threat report, the procedure examples are based on information extracted from reports and provide an aggregate view of the available information about that event from numerous sources. Each procedure example from these three events was reviewed to extract/synthesize and document the data or technical artifacts that can be used to detect the activity. The key criteria reviewed for each procedure example are defined below.
  • Data Source (1): Each technique within the ATT&CK knowledge base is mapped to a set of Data Sources that could potentially be used to detect that activity. [^7] https://attack.mitre.org/datasources/ This specifies what ATT&CK for ICS data source would have been most effective in detecting this procedure example.
  • Supporting Artifacts (2): This more precisely defines the technical artifact that would support the detection. This identifies the specific information that needs to be shared to enable the receiving party to detect the same threat/procedure example on their network. Supporting artifacts are not explicitly defined in ATT&CK, but they can often be derived from the procedure descriptions.
  • STIX Observable Supported (3): This evaluates whether the artifacts needed to detect that threat are documented as SCO. This is important because unless an artifact is documented by STIX, it cannot be formally communicated through STIX-based threat intelligence reports. These will be categorized into the following three types. While an artifact could have multiple possible STIX observable mappings, this example only illustrates one possible mapping that supports its associated categorization.
    • No Support: STIX lacks a relevant observable object that could correctly represent the identified artifact, thereby limiting the sharing of an observable.
    • Partial Support: STIX has an observable that could represent this artifact, but the observable does not fully encompass the semantics of the artifact, or the observable does not define significant fields to completely represent it. The partial nature implies ambiguity in the representation of observable artifacts, which in turn may yield unreliable detections or only support a limited set of aspects of the behavior. For example, a network traffic object may be used to represent an ICS command, with the protocol field representing the relevant ICS protocol (e.g., Supervisory Control and Data Acquisition (SCADA)) but missing specific support to represent the command itself.
    • Full Support: An existing STIX observable is fully capable of documenting this artifact, thereby enabling the creation of robust detection for the behavior.
  • Artifact Detailedness level (4): This specifies whether the information necessary to detect this specific threat behavior is available from the threat intelligence. This criterion is not intended to provide criticisms of the specific reports. Rather, it is intended to help identify whether information necessary to detect the threat is consistently available, or whether broader challenges exist in collecting the technical artifact that inhibit its sharing. For example, prior victims may be lacking adequate detection infrastructure or forensic capabilities to collect the evidence. These will be categorized into the following four types:
    • Missing: The procedure description does not mention the identified artifact nor any details about it.
    • Mentioned: The artifact was mentioned in the procedure description, but not addressed.
    • Described: Notable specific or distinguishing details about the artifact are described in the procedure description, but without sufficient unique information that would allow detecting it. These kinds of artifacts are non-searchable.
    • Actionable: The artifact is fully described in the procedure description with unique and specific information, in a way that a detection rule can be formed to match it with low false positives. These artifacts are searchable and can drive an automated response; they can be operated immediately or after a simple transform.
  • Proprietary artifacts (5): This identifies whether the artifact requires the understanding of some proprietary technology developed by a vendor/OEM, or whether that artifact is based on open/public standards or technologies. For example, many network protocols used by ICS devices lack public documentation of their functionality, while devices may contain proprietary logs or files with complex structures and content. Developing detections based on proprietary artifacts presents significant difficulty, as organizations often lack sufficient information on how to properly collect, parse, or analyze that data. These artifacts are categorized as follows:
    • Open/Standard Technology: These are technologies/protocols with publicly accessible documentation or are defined within an industry standard.
    • Proprietary - Documented Technology: These are also technologies/protocols developed and owned by a specific organization or company; however, in these cases, the documentation is publicly available. It is worth noting that such protocols often include undocumented features that may be relevant to the behavior of threat actors.
    • Proprietary - Undocumented Technology: These are technologies/protocols developed and owned by a specific organization or company. This refers to protocols for which documents are not available. The organization or company likely has internal documents, but they are not shared publicly. Proprietary but undocumented technologies can limit an organization’s ability to collect and analyze the data necessary to perform various security functions.
  • Protocol Parsers (6): If the artifact and associated data source include the detection of a network protocol, this identifies whether there are publicly available parsers for that protocol.
4.1 Deep Dive into Threat Information of Three Major Attacks
In this section, we review the procedure examples of Triton, Stuxnet, and Industroyer from the MITRE ATT&CK for ICS according to the criteria defined in the previous section.
4.1.1 Triton
The Triton malware [CISA2017HatMan], also known as TRISIS or HatMan, targeted safety instrumented systems within an oil refinery by manipulating both an engineering workstation and the Triconex safety controllers used in the sulfur recovery units and burn management systems. While the broader campaign during this attack included activities used to gain access to the facility (references as TEMP.Veles in ATT&CK), the defined Tactics, Techniques and Proceduress specific to the ICS environment primarily address the behaviors associated with the manipulation of Windows-based engineering workstations, safety controllers, and the associated network communications between them. The following sections will review the adversary activity associated with these three areas based on the previously defined criteria.
TriStation Workstation (Table 1): The Triton malware initially targeted the engineering workstation and associated TriStation software running on that device. The malware included a Python program that was compiled into an EXE, but which included a file name that masquerades as the authentic TriStation executable [T0849]. The malware then used the same Application Programming Interface (API) that TriStation uses to communicate with and manage the Triconex devices to access those devices [T0834]. Because the workstation is a traditional Windows-based platform, both of these techniques could be determined by reviewing files matching the hash of the EXE file, leveraging the STIX File: Hash object. This hash value was provided in the threat intelligence reports, and the Windows-based functionality is well-documented. However, detecting the malicious API uses likely depends on the runtime analysis of the specific system calls used by the TriStation workstations, which are not specified in any documentation, and there is no available STIX object to represent these system calls.
Table 1: Analysis of Triton Software TriStation Workstation Activity.
Tech. 1-Data Source 2-Artifact 3-STIX 4-Artifact Detailedness 5-Proprietary 6-Parser
T0849 File-Metadata File hash of exe Full - File:hashes Actionable Proprietary - Documented NA - host
T0853 Command/ Process File hash of exe Full - File:hashes Actionable
T0871 Process - Operating System (OS) API Exec TriStation API calls No Actionable Proprietary - Undocumented
TriStation Protocol (Table 2): The TriStation protocol, which is a proprietary protocol without any available open-source parsers, was heavily used throughout the attack. The adversarial techniques using this protocol include discovering target devices [T0846], detecting and changing their operating mode [T0858] [T0868], downloading programs to the targeted safety controller [T0843], executing the program [T0871], and extracting programs from the device [T0845]. Developing detections for this activity requires the ability to (i) understand and parse the TriStation protocol  [tristation_protocol_reversing], (ii) identify specific functions of the protocol, and (iii) represent function payloads (e.g., program downloads contents). The STIX Network Traffic observable could be used to represent this activity, specifically the src/dst_payload_ref objects, which reference specific strings within the protocol payloads. However, using the src/dst_payload_ref STIX objects for the TriStation protocols techniques, any detection utilizing these objects is likely to be fragile.
Table 2: Analysis of Triton Software TriStation Protocol Activity.
Tech. 1-Data Source 2-Artifact 3-STIX 4-Artifact Detailedness 5-Proprietary 6-Parser
T0846 Network Traffic Content Parsing of TriStation Protocol – Device discovery Partial – Network Traffic src/ dst_payload_ ref Described Proprietary - Undocumented No
T0858 - Operating mode changes
T0868 - Operating mode status
T0843 - Program downloads/appends
T0845 - Program uploads
T0885 Custom malware C2 payload Actionable NA NA
Triconex (Table 3): Triton’s targeting of the Triconex safety controller required multiple adversary techniques used to manipulate devices that execute custom logic. Once the malicious program was downloaded to the device, it was executed on the device [T0821]. Then, the malware gained excessive privileges by executing a vulnerable system call on the device [T0890], and manipulated the running firmware of the device [T0857] to maintain an implant. Next, it enabled Command and Control (C2) [T0869] by manipulating a protocol handler to link to malicious code injected into the device [T0874]. Detecting this activity requires the ability to extract and audit programs on the device, such as through program upload functions or integrity checks. This could potentially be detected by the Process: Hash STIX object, while many devices produce CRC’s or checksums of deployed programs, these could be easily converted to a hash to support this object type. These Cyclic Redundancy Checks within the Triconex devices are a proprietary but documented function [triconex_safety_evaluation]. Most of these other Triconex behaviors are difficult to detect because they manipulate the device’s run-time memory, which depends on the undocumented and proprietary operation of the device. Detecting the exploitation of a vulnerable system call typically requires a runtime agent running on the device, while detecting the runtime manipulation in the underlying firmware requires mechanisms that perform integrity checks of a device’s OS and runtime environments. Therefore, these specific techniques likely cannot be detected through traditional detection platforms and STIX based information sharing.
Table 3: Analysis of Triton Software Triconex Activity.
Tech. 1-Data Source 2-Artifact 3-STIX 4-Artifact Detailedness 5-Proprietary 6-Parser
T0821 Asset Software PLC program CRC Partial – Process: Hash Actionable Proprietary - Documented NA - host
T0890 Process - OS API Exec OS API calls No Described Proprietary - Undocumented
T0834
T0857 Firmware FW/Memory Contents Partial – Process: Hash
T0874
4.1.2 Stuxnet
The Stuxnet malware targeted Siemens PLCs used to support uranium enrichment cascades. The malware targeted both the PLCs, by manipulating the application program deployed on the device, and the engineering software deployed on the workstation (WinCC). Therefore, the associated TTPs then heavily focuses on the manipulation of the PLC, the software on the engineering workstation, and their associated network communications. Tables 4-6 enumerate each adversarial technique defined within ATT&CK for ICS associated with the Stuxnet malware.
SIMATIC WinCC Workstation (Table 4): The malware was initially deployed to engineering workstations using the SIMATIC WinCC software using either hard-coded credentials [T0891] or by copying itself to connected USB drives [T0847]. The hard-coded credential used could have been represented by the STIX User Account: Credential observable to enable sharing this information. The malware also propagates by modifying the content of SIMATIC Step 7888https://www.siemens.com/global/en/products/automation/industry-software/automation-software/tia-portal/software/step7-tia-portal.html project files found on the device by searching for files with the appropriate Step 7 format [T0873]. The modified project files could be detected by sharing the malicious contents or by representing the overall hashes of the activity, either through the STIX File: String or File: Hash observable.
Then, the malware installs a malicious dynamic-link library (DLL) which masquerades as an authentic file [T0849], the hash of this file could be represented by the File: Hash STIX observable. This DLL further injects any project files loaded [T0863]. The malicious DLL identifies targets by reviewing the contents of incoming data blocks to identify specific devices [T0888]. The DLL then downloads code to the PLC [T0843] and manipulates data sent between the PLC and workstation [T0851]. Since these activities all require understanding the execution of DLL, detecting this activity would require an understanding of specific function calls both within the DLL and OS system calls. However, as previously stated, STIX currently lacks support for this.
Table 4: Analysis of Stuxnet WinCC Workstation Activity.
Tech. 1-Data Source 2-Artifact 3-STIX 4-Artifact Detailedness 5-Proprietary 6-Parser
T0891 Logon Session - Metadata SQL Server password Full – User Acct: Cred Actionable Proprietary - Documented NA - host
T0847 File - Metadata Malicious file names Full – File: Name Actionable Proprietary - Documented NA - host
T0873 File - Metadata PLC Project Files Full – File: String/Hash Described Proprietary - Documented NA - host
T0849 File - Metadata DLL hash Full – File: String/Hash Actionable Proprietary - Documented NA - host
T0863 Command Execution / Process Creation Command name Full - File: Name Described Proprietary - Documented NA - host
T0874 Process – OS API Exec OS API calls No Mentioned Proprietary - Documented NA - host
T0888
T0851
Network Activity (Table 5): Stuxnet used a wide array of network activity to both propagate malware across devices and also to manipulate the control of the operational process. The malware utilized various SQL server-stored procedures to transfer files and execute malware [T0886], including transferring the Stuxnet DLL to other systems [T0867] and executing the malware [T0866]. This would require monitoring the SQL Server protocol for the execution of those specific stored procedures. This could be potentially shared by using the STIX Network Traffic: src/dst_payload_ref observable; however, it is unclear whether the SQL protocols could be effectively represented by payload strings. Then, the malware attempts to establish a C2 connection to external servers [T0885] over port 80, which could have been shared using the STIX Network Traffic: dst_port observable.
The Siemens S7Comm protocol was used to perform program downloads from the WinCC workstation to the PLCs [T0843]. Once the malicious code was deployed to the PLC, it then sends Profibus messages to frequency converter drives [T0836]. While Profibus is an International Electrotechnical Commission (IEC) standard, S7Comm (Siemens) is a proprietary protocol, though both have publicly available parsers. These activities would also likely need to be shared using the STIX Network Traffic: src/dst_payload_ref observable, though it is unclear whether these behaviors could be effectively shared using it.
Table 5: Analysis of Stuxnet Network Activity.
Tech. 1-Data Source 2-Artifact 3-STIX 4-Artifact Detailedness 5-Proprietary 6-Parser
T0866 Network Traffic Content SQL Strings/ Protocol Parser Partial – Network Traffic: src/ dst_payload_ ref Actionable Proprietary - Documented Yes
T0867
T0885 Network Traffic Connection Connection to Port 80 Full – Network Traffic: dst _port Mentioned Proprietary - Documented NA
T0843 Network Traffic Content Parsing of S7Comm Protocol Functions: Program Downloads/Appends Partial – Network Traffic: src/ dst_payload_ ref Missing Proprietary - Undocumented Yes
T0836 Network Traffic Content Profibus Network Traffic Partial – Network Traffic: src/ dst_payload_ ref Described Open / Standard No
Siemens PLCs (Table 6): The malware deploys malicious programs by manipulating controller tasking (e.g., OB1, OB35) [T0851][T0821] on the PLCs. The manipulation of controller tasking and programs can be detected by either performing program uploads or performing integrity checks. These could be partially represented by File: Hash STIX observables, since many PLC application programs are frequently represented by other message digest mechanisms, such as PLCs or checksums. This malicious program then performs a number of additional techniques on the PLCs, including sniffing Profibus message contents [T0842] and monitoring the state of the operational process [T0801] and calls an existing system function block used by the device [T0834]. The program collects data from the I/O image [T0877] and then intercept/overwrite peripheral output to prevent detection [T0835]. These could either be detected through some sort of run-time analysis, by verifying the specific OS API calls made, or by extracting and performing an analysis of the malicious program performed on the device. The former would depend on the device’s ability to perform run-time monitoring of device functions; this capability is typically not supported by PLCs. Since most PLCs do support application program uploads, the specific PLC API calls can be analyzed to observe this behavior. Unfortunately, this presents a clear gap in STIX, as there is currently no way to reference specific PLC application logic functions.
Table 6: Analysis of Stuxnet PLC Activity.
Tech. 1-Data Source 2-Artifact 3-STIX 4-Artifact Detailedness 5-Proprietary 6-Parser
T0821 Asset – Software PLC code Partial – File: Hash Described Proprietary - Documented NA - Host
T0842 Asset – Software / Process – OS API Exec PLC OS API calls No Described Proprietary - Documented NA - Host
T0801 Asset – Software / Process – OS API Exec PLC OS API calls
T0869 Asset – Software / Process – OS API Exec PLC OS API calls Actionable
T0834 Asset – Software / Process – OS API Exec PLC OS API calls Described
T0877 Asset – Software / Process – OS API Exec PLC OS API calls Actionable
T0835 Asset – Software / Process – OS API Exec PLC OS API calls Described
4.1.3 Industroyer
The Industroyer or CrashOverride malware was used as part of an attack that occurred in December 2016, targeting Ukraine’s power grid and causing a blackout in the northern part of the capital, Kyiv. It was capable of operating a substation’s switches and circuit breakers, utilizing several common industrial communication protocols, including IEC 60870-5-101, IEC 60870-5-104, IEC 61850, and OPC DA. Additionally, this malware features a built-in "wiper" module that can erase the configuration files for the equipment controlling the circuit breakers, making recovery more challenging.
Windows Activity (Table 7): The malware launcher is initially executed as a .dll through a command line parameter [T0807], and multiple other modules were also executed as .dlls. This could be represented as a STIX File Name object, while the initial launcher .dll name is not included in the report, the associated wiper .dlls are provided. The malware then enumerates systems call on the system, which requires executing Windows system calls [T0840]. Furthermore, the malware connects to COM ports on the Windows system, which involves executing Windows system calls. However, neither of these system-call-focused artifacts can be represented by STIX observables [T0803] [T0804] [T0805]. Later, the malware executes a wiper that disables system services and wipes files on the device. While the specific system calls used to perform this are not defined, there are likely a limited set of Windows calls that enable this. Furthermore, STIX lacks an observable to represent system calls used to understand runtime malware functions.
Table 7: Analysis of Industroyer Windows Activity.
Tech. 1-Data Source 2-Artifact 3-STIX 4-Artifact Detailedness 5-Proprietary 6-Parser
T0807 Command Execution CMD line parameter w/ DLL Full – File: Name Actionable Proprietary - Documented NA - Host
T0881 Service - Metadata / Process – OS API Exec OS API call associated with service stop No Actionable Proprietary - Documented NA - Host
T0809 Process – OS API Exec OS API calls associated with wiper functions No Actionable Proprietary - Documented NA - Host
T0840 Process – OS API Exec OS API calls associated with network enumeration No Actionable Proprietary - Documented NA - Host
T0803 Process – OS API Exec System Calls for COM Ports No Mentioned Proprietary - Documented NA - Host
T0804
T0805
Network Activity (Table 8): The first network-focused activity is the establishment of a proxy through port 3128 [T0884], which can be represented by the STIX Network Traffic src/dest_port observable. Then, the malware has a broad set of behaviors that utilize a set of automation protocols to interact with substation devices, including supporting remote discovery, collection, monitoring, and sending unauthorized command messages.
At least two protocols, OPC and 61850 leveraged discover functions within the protocols to identify devices supporting those protocols [T0846]. Then, several protocol functions were utilized to collect and monitor the status of devices using the 61850, OPC, and IEC 104 protocols [T0888][T0801]. The unauthorized command messages were used to manipulate the operation of devices, specifically to open/close breakers [T0855][T0806]. While none of these protocols have specific representation in STIX, they could be partially represented by the Net Traffic src/dst_payload_ref object. Fortunately, these are all open or standards-based protocols, and all have available parsers.
Table 8: Analysis of Industroyer Network Activity.
Tech. 1-Data Source 2-Artifact 3-STIX 4-Artifact Detailedness 5-Proprietary 6-Parser
T0884 Network Traffic Connection Connection to Port 3128 Full – Network Traffic: src/dest_port Described Open / Standard NA
T0846 Network Traffic Content 61850 discovery functions Partial Network Traffic: src/ dst_payload_ ref Described Open / Standard Yes
T0846 Network Traffic Content OPC discovery functions Partial Network Traffic: src/ dst_payload_ ref Actionable Open / Standard Yes
T0801 OPC/61850 monitor functions
T0888 61850 collection functions
T0888 OPC collection functions
T0888 104 collection functions
T0855 104 command functions
T0806 104 command functions
SIPROTEC/Relay Activity (Table 9): Industroyer also has the capability to use the CVE-2015-5374 vulnerability to make a SIPROTEC relay unresponsive. The malware sends a SIPROTECH message to a relay, causing the device to enter an unresponsive state, requiring the device to be rebooted [T0814][T0816]. This message is an 18-byte packet to a proprietary UDP port 50,000 on the device. This 18-byte string is provided in the threat intelligence and could be represented by the STIX payload_ref object; however, it is unclear whether this is a reliable or precise indicator, as it is a proprietary protocol. On the device side, the message initiates an incomplete firmware update, which is not completed, leaving the device unresponsive [T0800]. This is potentially available in device application logs; however, there is no detail about this activity from the device side, and no existing STIX observable to support log events.
Table 9: Analysis of Industroyer Network Activity.
Tech. 1-Data Source 2-Artifact 3-STIX 4-Artifact Detailedness 5-Proprietary 6-Parser
T0814 Network Traffic Content Strings/ SIPROTECH protocol parser & mgmt functions Partial – Network Traffic: src/ dst_payload_ ref Actionable Proprietary - Undocumented No
T0816
T0800 Application Log Content Device Log Events No Missing Proprietary - Undocumented NA - host
4.2 Thorough Analysis of Threat Information Sharing at ATT&CK for ICS
4.2.1 Methodology
To provide a comprehensive assessment beyond the three detailed case studies, we conducted a systematic analysis of all 196 procedure examples across 79 techniques documented in ATT&CK for ICS. Our methodology consists of four phases:
  1. 1.
    Data Acquisition and Preparation - All procedure descriptions were retrieved from the public MITRE ATT&CK STIX data for ICS [^9] https://github.com/mitre-attack/attack-stix-data/blob/master/ics-attack/ics-attack.json and normalized into plain-text records. Each record retained its corresponding technique identifier, and malware family linkage.
  2. 2.
    Automated Observable Extraction - Each procedure description was processed by a LLM (GPT o3-mini) using a structured prompt that explicitly defined the task, decision rules, and expected JSON output format. The model was instructed to identify every observable mentioned in the text, ranging from file names and commands to ICS-specific protocol elements, and to classify each one according to the pre-defined attributes.
  3. 3.
    Observable Categorization - Every extracted observable was represented as a structured schema record containing the key criteria we defined before.
  4. 4.
    Post-processing and Quality Control - According to the STIX 2.1 specification, the software Cyber-Observable Object is defined as "high-level properties associated with software, including software products". [^10] https://docs.oasis-open.org/cti/stix/v2.1/csprd01/stix-v2.1-csprd01.html#_Toc16070740 This equivocal definition can mislead an LLM into labelling malware identifiers as observables. Yet, STIX classifies malware itself as a distinct Domain Object [^11] https://docs.oasis-open.org/cti/stix/v2.1/csprd01/stix-v2.1-csprd01.html#_Toc16070566 rather than as an observable artifact. Because SCOs are intended to capture low-level forensic evidence, while Domain Objects (e.g., malware encodes higher-order threat-intelligence concepts), keeping malware entries in the observable layer would blur these semantic boundaries. [^12] https://docs.oasis-open.org/cti/stix/v2.1/csprd01/stix-v2.1-csprd01.html#_Toc16070567 To preserve the intended separation of concerns, we therefore filtered out every item that the model had incorrectly flagged as malware-related. The resulting dataset was manually reviewed to eliminate duplicates, to remove entities incorrectly labeled as Observables, and to correct misclassifications. Overall, out of 361 observables, 103 observables’ data were manually modified:
    • 34 observables specifications were modified to be more ICS specific domain
    • 27 observables were manually classified as Actionable instead of Described
[…]
The systematic analysis revealed patterns that directly address our research questions. The following sections present the empirical findings, organized by: (i) STIX representation gaps; (ii) actionability of extracted observables; and (iii) dependency on proprietary technologies.
4.2.2 Observation and Analysis
STIX GAP Taxonomy
Out of 361 extracted observables, over half of the observables (191, around 53%) have only partial support in STIX 2.1, meaning they can only be approximated via generic or custom objects. About 19% (69 observables) have no support at all in STIX (i.e., no relevant cyber observable exists), while only around 28% (101 observables) are fully supported by an existing STIX object type. This indicates a significant coverage gap for many ICS-specific artifact types.
  1. 1.
    Fully Supported (101) - These are mostly standard Information Technology (IT) indicators (e.g., files, network addresses, domain names) which map cleanly to STIX SCOs like File, IPv4-Addr, Domain-Name, etc. Few ICS-specific items fall in this category, underscoring that STIX 2.1’s built-in types were not designed with ICS nuances in mind.
  2. 2.
    Partially Supported (191) - The majority of ICS observables fall here. They often correspond to ICS domain concepts that lack a first-class STIX object, but can be accommodated within a generic object or a custom extension. For example: MMSgetNameList (an ICS protocol command) is labeled Partial: network-traffic, meaning it can be captured as network traffic data, but STIX has no explicit "ICS command" type. Many PLC programming functions (e.g., Siemens S7 functions like SFC1 or block names like OB1) were marked Partial: artifact, indicating they could only be represented as generic file/artifact content.
  3. 3.
    Not Supported (69) - These are observables with no meaningful STIX representation. They include highly ICS-specific elements for which even an approximation is lacking. For example: ctlSelOn – an ICS data tag/field from a control system historian – is marked "No support" because STIX has no concept of a tag or point name in a PLC/SCADA database. Likewise, ICS memory addresses and register values (e.g., Information Object Addressess in IEC-104) are also unsupported. In general, ICS data tags, device configurations, and field values (e.g., status flags, function codes) show up repeatedly among the "No support" entries. This trend highlights that industrial control system observables are a major gap area in STIX 2.1, the schema covers IT-centric artifacts well, but things like control logic variables, PLC memory blocks, or proprietary protocol fields fall outside its current scope.
Trends in Unsupported Types.
The unsupported entries tend to cluster in specific classifications. Many ICS Data Tag artifacts (control system historian points, device tags) had no STIX support. Even some general IT items like specific software tool names or report identifiers were marked unsupported when they do not correspond to STIX’s defined objects (e.g., an internal report ID or a code function name has no direct STIX analog). In contrast, common network or host indicators, such as files, IP addresses, and user accounts, are typically fully covered. This suggests that STIX 2.1’s coverage gap is most pronounced for specialized ICS domain artifacts, such as data related to industrial devices, proprietary protocols, and control logic, whereas typical enterprise observables are well-supported.
Artifact Detail Level Distribution
Figure 2: Distribution of Artifact Details within Top 10 Classification Categories ("Actionable" in blue, "Described" in teal, "Mentioned" in orange).
The observables were categorized into over 100 classification types, but a few dominate (see Figure 2). It is worth noting that an ICS data artifact class, ICS Data Tag (8), also makes the top 10, indicating that specific tag names/variables from control system databases were noted multiple times despite STIX gaps. In general, the classification distribution highlights that ICS/OT entities (devices, protocols, and commands) were prominent, alongside a mix of traditional IT indicators (files and network configurations) present in the threat reports.
Figure 3: Distribution of Artifact Details within Top 5 Data Sources("Actionable" in blue, "Described" in teal, "Mentioned" in orange).
Each observable was also tagged with a data source indicating where that artifact could be collected. The most common source by far is network traffic (93 observables). This aligns with the many ICS protocol messages and network indicators extracted (since those would be seen in network captures or Packet Captures). Notably, ICS historian logs (50) are the third-largest category; many of the ICS-specific observables, such as tag values and device status changes, are recorded in historian databases or engineering logs, emphasizing the importance of historian data in detecting ICS attacks. It is worth highlighting that ICS-specific telemetry (Historian, PLC ladder logic captures) together make up a significant chunk, demonstrating that to fully monitor these observables, one must pull from ICSs and not just traditional IT logs.
Actionable Observables Lacking STIX Support
One important observation is that several observables deemed "Actionable" (high-value IOCs) are not fully supported in STIX 2.1, which could hinder information sharing. In our results, 48 actionable observables (over half of all actionable ones) were only partial or not supported in STIX. These include ICS-specific items, such as ctlSelOn, which is a precise, unique tag name that an attacker can manipulate (actionable for detection in ICS logs); yet, STIX has no way to represent it except as a generic artifact or custom field. The same goes for certain PLC function codes or protocol commands that are very specific (and thus actionable within that context) but have no STIX presentation.
The key insight is that being "Actionable" does not guarantee STIX support. There is a misalignment where CTI producers can identify a highly specific ICS indicator, but when codifying it in STIX format to share with others, they struggle due to the schema’s limitations. This can complicate the intelligence sharing process for ICS threats; either the observable has to be dropped (if there’s no representation) or shared in an unstructured way (in notes or description fields), reducing its utility. It reinforces the earlier point that extending STIX or developing custom SCOs for ICS could significantly enhance our ability to share and leverage these crucial observables. The fact that dozens of actionable ICS indicators fall through the cracks in STIX 2.1 is a call to action for the community to address these schema gaps for industrial security.
5 Analysis of Vulnerability Information Sharing
In addition to studying cyber threat activity, the analysis also reviews the sharing of vulnerability advisories associated with known exploitable ICS vulnerabilities based on their inclusion in the CISA KEV Catalog. This list was produced by reviewing the KEV Catalog for Common Vulnerabilities and Exposuress identified since 2023, which are associated with CISA ICS Alerts. Since these vulnerabilities have been known to be exploited by threat actors, they present a clear risk to asset owner environments, requiring some mitigating actions. While within IT environments these vulnerabilities are typically patched, it is well documented asset owners struggle to quickly deploy patches [sans_ics_ot_cybersecurity]. If an asset owner cannot quickly patch these vulnerabilities to mitigate their risk, they may need to deploy detection mechanisms to prevent their exploitation. Therefore, vulnerability information is reviewed in accordance with the same requirements as defined in Section 2. That is, evaluating whether the vendor provided some rule to support the detection of the vulnerability, or whether sufficient technical details are available about the product and associated vulnerability for the asset owner to generate their own detection.
Table 10 documents the analysis of the 9 CISA vulnerability advisories and the associated vendor-provided vulnerability advisories. For each advisory, the CISA ICS Advisory number, the associated CVEs, vendor, and associated protocols (if applicable) are defined. Each advisory was reviewed for the following criteria.
  1. 1.
    Rules/Analytics Available (1): This identifies whether the advisory provides any information about detection rules (e.g., a Snort rule) or analytics that can be deployed to detect the advisories.
  2. 2.
    Availability of Technical Details (2): This criterion assesses whether the report or advisory includes sufficient technical information such that a detection rule or analytic could be developed based on that information. While many vulnerability advisories provide high-level discussion about the vulnerability and mappings to known Common Weakness Enumerations, they also do not provide the technical details required to develop rules or analytics to detect exploitation. If technical information is available but insufficient to inform a detection, it is labeled as ’Partial.’
  3. 3.
    Proprietary (3), and Parsers (4): For vulnerabilities that are associated with a network protocol, this explores whether the protocol is open, proprietary, and documented, or proprietary and undocumented (3). Next, it reviews whether a parser exist for that protocol (4). If a parser exists but does not fully cover the protocol’s use or implementation, it is labeled as Partial.
Table 10: Table 10: Review of 9 ICS vulnerabilities with known exploitation.
ICSA Alert CVE Vendor Protocol 1- Rules 2 - Tech Detail 3 - Proprietary 4 - Parsers
23-355-01 2023-49897 FXC HTTP No No Open - HTTP Partial
23-355-02 2023-47565 QNAP HTTP No No Open - HTTP Partial
23-348-15
2023-4911
2023-44487
Siemens NA - Host No Partial N/A N/A
23-320-11 2023-6448 Unitronics PCOM No No
Proprietary -
Undocumented
No
23-297-01
2023-20273
2023-20198
Rockwell
Automation
HTTP
Yes -
External
Partial Open - HTTP Partial
23-264-05 2020-16017
Rockwell
Automation
HTTP No Partial Open - HTTP Partial
23-075-01 2021-4034 Siemens NA - Host No Yes N/A N/A
23-193-01
Documented
exploit but
not in KEV
2023-3595
2023-3596
Rockwell
Automation
CIP Yes Yes
Proprietary -
Documented
Yes
5.1 Key Outcomes
The outcome of this analysis highlights that current vulnerability advisories lack sufficient technical details for asset owners to create their own detections. Only two out of the nine vendor advisories reviewed include a detection rule or analytic for the vulnerabilities. Furthermore, vulnerability advisories often lack the necessary technical details to support the development of effective detections. Four provided no technical details, excluding the two vulnerabilities that included detection rules, while three provided partial technical discussions, which were insufficient for developing a detection. Further, one protocol was fully proprietary and undocumented. At the same time, five were based on HTTP, an open protocol, but can have highly diverse specific customization, making parsing difficult without details of the specific application functions. While parsers for HTTP exist, they would still need to be heavily tailored towards the particular device’s application. This analysis highlights that detection rules are not consistently provided and critical information about device vulnerabilities or associated protocols not provided with sufficient detail to enable the development of detections. Based on this analysis, it’s unclear how asset owners are expected to protect themselves from the reported vulnerabilities.
6 Challenges in Threat Information Sharing
The prior sections provide an evidence-driven review of both available threat intelligence data (through ATT&CK for ICS), and recent vulnerabilities advisories defined by vendors and CISA. This analysis identifies four key challenges that should be addressed to ensure threat information can be more shared more effectively to ensure asset owners can effectively detect these threats.
  1. 1.
    A lack of defined technical artifacts to support sharing of indicators of OT threats through STIX. This analysis demonstrated that while existing STIX Observables would be sufficient for detecting a large majority of the TTPs identified in the previous analysis, there remain multiple TTPs that lack sufficiently defined technical artifacts. Therefore, expanding STIX observables is necessary to support the sharing of all known adversarial techniques associated with ICS incidents. Key examples include:
    • PLC Code/Logic: Numerous techniques include the deployment of malicious logic on a PLC. While many PLCs leverage vendor-specific tailoring of their logic, many devices share some fundamental capabilities and languages, especially as they align with the IEC 61131-3 standard for developing programs. These programs are typically compiled before deployment to the controller. However, the source code is typically transferred alongside the binaries to support diagnostics of the deployed program. Therefore, mechanisms to support understanding program segments on the controller, both through source and compiled code, may be effective in detecting malicious logic.
    • Integrity checks: Further, many controllers also use integrity check mechanisms, typically based on CRCs or Checksums, to verify that programs have not changed  [plc_security_practices]. On the one hand, security products can verify that a specific program has been downloaded to the device by validating its signatures. On the other hand, CRC/Checksum integrity checks are not sufficient against threat actors who intentionally modify the PLC code  [schuett2014evaluation]. There is a need for a wider deployment of secured integrity signatures  [banerjee2024designing].
    • ICS Protocols: Currently, the STIX standard supports prevalent IT protocols, including HTTP and SMTP. However, each incident reviewed above has techniques that utilize ICS-specific protocols. While these could potentially be parsed through more general NetworkTraffic payload objects, the associated observables will still likely be ineffective without tailored objects. Therefore, developing observable objects to support key ICS protocols and fields would improve the effectiveness of sharing and detection of techniques targeting those protocols. It may be possible to craft a general ICS protocol object using typical functionality associated with ICS protocols.
  2. 2.
    The dependence on proprietary technologies and protocols without available parsers hinders the ability to develop detections. While vendor technologies may require custom/proprietary protocols or the customization of standard protocols to address unique challenges faced when interfacing with devices, these protocols are increasingly targeted by threat actors, and therefore necessitate the ability to be monitored for threats by the asset owner. The following table reviews the unique protocols identified in this analysis. Of the nine overall protocols, at least three proprietary protocols were identified, and two had no publicly available parsers. Unfortunately, organizations are required to monitor network traffic in order to identify the behavior of threat actors. However, this dependency on proprietary and undocumented protocols, along with limited access to parsers, prevents the deployment of the necessary detection capabilities.
    Table 11: Available Network Parsers for ICS Protocols.
    Threat/vuln Protocol Parsers available
    Industoryer IEC 61850 MMS Yes
    IEC 60870-5-104 Yes
    OPC DA No
    Stuxnet S7Comm Yes
    MS SQL(TDS) Yes
    Profibus No, but Profinet does
    Triton Tristation No
    Icsa-23-348-15 PCOM No
    Icsa-23-193-01 CIP Yes
    While proprietary and undocumented protocols were acceptable when vendors could assume only their own devices would operate within a closed ecosystem, the need to integrate network monitoring capabilities invalidates this assumption. Therefore, vendors must provide standard documentation and available parsers for network protocols to ensure asset owners can sufficiently detect threat actor behaviors in their environments.
  3. 3.
    Vulnerability advisories lack sufficient technical details to develop a detection of exploited vulnerability and lack rules/analytics to support that detection. Current industry norms suggest that vendors are responsible for providing timely patches or updates to address product vulnerabilities, thereby mitigating the associated risks. However, it is well known that most asset owners struggle to perform patching in OT environments due to shortages of OT security personnel, limited maintenance periods, and high availability requirements. If a patch cannot be deployed, the organization may not have an alternative mechanism to mitigate that risk. It may only be able to monitor for any exploitation activity of this vulnerability. Unfortunately, vulnerability advisories constantly lacked an associated detection rule or analytic, and also lacked sufficient technical details necessary to devise an effective rule. Specifically, in the sampling of the nine vendor advisories reviewed for this study, five did not have mitigations other than patching, and seven lacked sufficient detail to support the detection, and only two included actual detection rules. Without sufficient technical details about vulnerabilities, asset owners cannot effectively develop their own detection rules or analytics, as this would likely require time- and resource-intensive reverse engineering of the device and associated protocols. Organizations are therefore primarily dependent on external security firms to perform these actions, but this ultimately presents additional costs to the organizations, both through the cost of the detection product and through the staff time dealing with false positives from an ineffective or imprecise rule. The two vulnerability advisories that contribute to associated detection rules demonstrate the feasibility of publishing these rules and should be acknowledged as an exemplar of leading industry practices.
  4. 4.
    Threat reports/advisories lack sufficient technical details or supporting technical artifacts across many procedure examples, constraining the development of detection rules or analytics. Similar to the above challenge regarding the lack of technical details about vulnerabilities and associated detection rules or alerts, threat advisories do not consistently provide sufficient technical details or analytics/rules to enable the development of detections. In the sampling of the 51 procedure examples identified for this study, at least 44 did not have sufficient technical detail to create a detection. However, unlike vulnerability advisories, security firms or government agencies typically release threat advisories. These organizations typically have limited access to information about threat activity due to constraints on the available data sources required for forensically analyzing the environments and devices. Furthermore, as previously identified, detection vendors typically have limited information about vendor technologies, which may require additional reverse engineering to fully understand the threats’ technical actions and develop the required detections. Therefore, this challenge should not be perceived as an inadequacy or deficiency of the organizations developing these reports, especially when the developing organization is producing this information to support their commercial product. Instead, this is only intended to highlight the need for a more consistent approach/format for the structure and content of threat advisories. While many reports document specific ATT&CK TTPs, they lack specific technical artifacts or technical details necessary to detect that threat.
7 Limitations and Implications
This study, while comprehensive in scope, has several limitations that should be considered when interpreting the findings. The first limitation is that our analysis primarily relies on MITRE ATT&CK for ICS as the aggregated source of threat intelligence. While ATT&CK represents one of the most comprehensive publicly available knowledge bases of ICS threats, this choice introduces potential coverage bias. ATT&CK may overrepresent well-documented, publicly disclosed incidents while underrepresenting threats that remain classified or proprietary. If alternative sources such as sector-specific Information Sharing and Analysis Centers, regional CERTs, or commercial threat intelligence were analyzed, we might observe different patterns in observable types and documentation quality. A second important limitation arises from the methodological constraints. While we validated the LLM-based extraction through manual review and correction of 103 observables, automated extraction may have missed implicit observables or misclassified complex ICS-specific artifacts. The 87% inter-rater reliability, while acceptable, indicates some subjective interpretation in observable classification.
8 Conclusion
This study presents the first comprehensive, evidence-based analysis of the challenges associated with threat information sharing in industrial control systems. Through a deep analysis of threat information from three major attacks, a systematic examination of 196 procedure examples from ATT&CK for ICS, and nine recent vulnerability advisories, we have empirically demonstrated four critical barriers that impede effective threat intelligence sharing in OT environments.
Our findings reveal a fundamental misalignment between the information needed for ICS threat detection and the capabilities of current sharing mechanisms. The analysis reveals that 53% of ICS-specific observables lack adequate STIX representation, and 19% have no STIX representation, while 77% of extracted artifacts lack sufficient technical detail for operational use. These are not merely technical oversights but systemic issues stemming from the unique characteristics of OT environment’s proprietary protocols, specialized devices, and domain-specific operational constraints.
The implications extend beyond technical standards. Our results suggest that effective ICS threat intelligence sharing requires a paradigm shift in how the community approaches information collection, representation, and dissemination. Vendors must recognize that security-by-obscurity, achieved through undocumented protocols, is no longer tenable when these same protocols become attack vectors. Security researchers must balance the need for operational security with providing actionable technical details. Standard bodies must evolve frameworks like STIX to accommodate ICS-specific artifacts that do not fit traditional IT paradigms.
Looking forward, this research establishes a baseline for measuring progress in ICS threat intelligence sharing. The methodology developed, combining manual analysis with automated extraction, provides a reproducible framework for ongoing assessment. As the ICS threat landscape continues to evolve, regular application of this methodology can track improvements in information sharing quality and identify emerging gaps.
The path forward requires coordinated action across multiple stakeholders. We recommend: (i) expansion of STIX to include ICS-specific observable types identified in this study; (ii) industry adoption of standardized vulnerability disclosure templates that include detection rules; (iii) development of open-source protocol parsers for critical ICS protocols; and (iv) establishment of sector-specific information sharing practices that balance operational security with actionable intelligence.
9 Summary
This paper presents a data-driven analysis of threat information sharing challenges in industrial control systems, examining real-world threat intelligence and vulnerability advisories. We analyzed 196 procedure examples across 79 techniques from the MITRE ATT&CK for ICS framework and nine recent CISA vulnerability advisories, identifying four key challenges:
  1. 1.
    Insufficient STIX Observable Support: 72% of ICS-specific threat observables cannot be adequately represented in current STIX standards, with critical gaps in PLC logic representation, ICS protocol specifics, and device-level artifacts.
  2. 2.
    Dependency on Proprietary Technologies: Critical threat detection depends on understanding undocumented proprietary protocols, with 4 out of 9 identified ICS protocols lacking publicly available parsers.
  3. 3.
    Inadequate Vulnerability Advisory Details:Only 22% of reviewed advisories provided detection rules, and 78% lacked sufficient technical detail for developing custom detections.
  4. 4.
    Limited Actionable Threat Intelligence: Only 23% of threat procedure examples contained actionable technical details, constraining defenders’ ability to develop effective detections.
Our systematic methodology, which combines manual analysis with automated LLM-based extraction, provides a reproducible framework for assessing the quality of threat intelligence. The findings demonstrate that current information sharing practices, while improved from previous years, still fail to meet the operational needs of ICS defenders. Addressing these challenges requires coordinated efforts from vendors, researchers, and standards bodies to develop ICS-specific sharing mechanisms that balance security concerns with operational utility.
APPENDIX
Observable Extraction Prompt
The following prompt was used to systematically extract observables from ATT&CK for ICS procedure descriptions using GPT o3-mini:
System Prompt: You are a helpful Cybersecurity assistant for identifying observables in Cyber Threat Intelligence text snippets. Task 1. You will receive a text snippet of a CTI report from a user. 2. Read the given snippet (plain text) carefully. 3. Extract every observable (artifact) mentioned – do not omit any. 4. For code snippets, include the full code, including triple backticks. 5. For each observable, output a JSON object with the exact fields listed in the Response format section. Definitions 1. Actionable Observable Unique & specific $\rightarrow$ a deterministic IDS/YARA/SIEM rule could match it with low FP. Immediately operable as-is (code snippet, exact URL, command, file name or path, API function) or after simple transform (e.g., Base64 decode, hash lookup, parameter substitution, memory dump). Searchable and can drive automated response. If the observable meets these criteria $\rightarrow$ artifact_details = "Actionable". 2. Described Observable Has notable specifics, but still not unique enough for detection; non-searchable. If the observable meets this criteria and not the Actionable Observable criteria $\rightarrow$ artifact_details = "Described". 3. Mentioned Observable A non-searchable observable, which doesn’t stand in the Actionable Observable criteria, nor the Described Observable criteria. For such observables $\rightarrow$ artifact_details = "Mentioned". 4. STIX Supported This evaluates whether the observable is documented as STIX 2.1 Cyber-observable Object Full: the observable’s type exists in STIX Cyber-Observable Objects. Partial: the observable does not map cleanly to a first-class STIX SCO, but can be approximated or expressed indirectly, or supported only via x_ custom properties or the generic artifact object. No: the observable isn’t Fully STIX supported nor Partially supported 5. Proprietary Artifact Open/Standard Technology Proprietary-Documented Technology Proprietary-Undocumented Technology Fields to produce for every observable Field Description observable_value Exact string (or faithful paraphrase). Escape any internal backticks. artifact_details "Mentioned" | "Described" | "Actionable" based on the definitions above. data_source Where it can be observed or collected (see cheat-sheet below). classification Short type label (e.g., "ICS Command", "URL", "Software/Tool"). STIX_supported "Full: <STIX_Object_Name>" | "Partial: <STIX_Object_Name>" | "No". proprietary_artifact "Open/Standard Technology" | "Proprietary-Documented Technology" | "Proprietary-Undocumented Technology". parser Known open-source/commercial parser name(s) for the data format, else null or "N/A" if not applicable. notes Any extra comments or context (Markdown allowed), or null if none. Common data_source cheat-sheet Network traffic • Netflow • PCAP • DNS logs • Web proxy logs • Endpoint (EDR) logs • System logs (Windows Event, syslog) • ICS historian • PLC ladder logic • Firewall logs • Cloud API audit logs • Memory dump • None (if not observable via telemetry) Response format (return only this JSON) 1{ 2 "observables": [ 3 { 4 "observable_value": "<VAL>", 5 "artifact_details": "Mentioned | Described | Actionable", 6 "data_source": "<text>", 7 "classification": "<one of allowed values>", 8 "STIX_supported": "Full: <STIX_Object_Name> | Partial: <STIX_Object_Name> | No", 9 "proprietary_artifact": "Open/Standard Technology | Proprietary-Documented Technology | Proprietary-Undocumented Technology", 10 "parser": "<text>" | null | "N/A", 11 "notes": "<text>" | null 12 } 13 ] 14}

### v2

@@
 As critical infrastructure increasingly becomes the target of cyber threats, organizations and governments must be able to detect and share information about threat actor activity.
-Numerous government reports, such as Executive Order 13691 [exec_order_13691] and the Cyberspace Solarium Commission report [cyberspace_solarium_2023], have highlighted the need for enhanced information sharing to safeguard critical infrastructure.
+Numerous government reports, such as Executive Order 13691 [14] and the Cyberspace Solarium Commission report [13], have highlighted the need for enhanced information sharing to safeguard critical infrastructure.
 Additionally, numerous government programs have been established to facilitate information sharing between the public and private sectors.
@@
 In this paper, we examine the challenges that hinder the sharing of actionable threat information with asset owners within the ICS domain.
-This study leverages the MITRE ATT&CK for ICS [mitre_attack_ics] knowledge base, which aggregates cyber threat intelligence reports across known ICS incidents.
-Our paper highlights the key challenges by reviewing three well-known incidents: Stuxnet [stuxnet_facts_report], Industroyer [ukraine_power_grid_analysis], and Triton [saudi_arabia_cyberattack].
+This study leverages the MITRE ATT&CK for ICS [36] knowledge base, which aggregates cyber threat intelligence reports across known ICS incidents.
+Our paper highlights the key challenges by reviewing three well-known incidents: Stuxnet [23], Industroyer [8], and Triton [29].
 This analysis aims to identify the adversarial techniques employed by threat actors and to determine the information that asset owners need to implement effective detection mechanisms against these techniques.
-We also explore nine recent ICS vulnerability advisories with evidence of exploitation as documented in the CISA Known Known Exploitable Vulnerability (KEV) Catalog  [cisa_kev_catalog].
+We also explore nine recent ICS vulnerability advisories with evidence of exploitation as documented in the CISA Known Known Exploitable Vulnerability (KEV) Catalog  [12].
 To provide a comprehensive assessment beyond these three high-profile incidents, this study extends the analysis to a systematic examination of 196 procedure examples across 79 techniques from 22 malware families documented in the MITRE ATT&CK for ICS framework.
@@
 Krasznay et al.
-[krasznay2021possibilities] highlighted the critical vulnerabilities in these systems, which stem from their original design priorities focused on safety rather than cybersecurity.
+[18] highlighted the critical vulnerabilities in these systems, which stem from their original design priorities focused on safety rather than cybersecurity.
 The increasing connectivity of ICS and Operational Technology (OT) systems has exposed these vulnerabilities to sophisticated adversaries, including state-sponsored actors, necessitating a paradigm shift in cybersecurity practices.
@@
 López-Morales et al.
-[lopez2024securing] extended these contributions by developing innovative methodologies that advance the state-of-the-art in CTI for CPSs.
-Their work introduces the ICS2 Matrix, an enhanced version of the MITRE ATT&CK framework tailored specifically to ICSs.
-The ICS2 Matrix incorporates novel attack techniques and mitigation strategies derived from an extensive analysis of Programmable Logic Controller (PLC)-related security research spanning 17 years.
+[19] extended these contributions by developing innovative methodologies that advance the state-of-the-art in CTI for CPSs.
+Their work introduces the $ICS^{2}$ Matrix, an enhanced version of the MITRE ATT&CK framework tailored specifically to ICSs.
+The $ICS^{2}$ Matrix incorporates novel attack techniques and mitigation strategies derived from an extensive analysis of Programmable Logic Controller (PLC)-related security research spanning 17 years.
 Additionally, the development of a high-interaction satellite honeypot provides a groundbreaking tool for simulating realistic adversary interactions, collecting data on emerging threats, and advancing the understanding of satellite-specific attack vectors.
 This work further proposes the creation of a sandbox environment for connected autonomous vehicles, enabling the simulation and analysis of cyberattacks on autonomous systems to improve resilience.
+Recently, MITRE extended its $D3FEND^{TM}$[22] cybersecurity ontology to OT, creating a structured knowledge base for defending cyber-physical systems, including new artifacts, such as controllers, sensors, and physical process actuators.
+Yet, their additional new artifacts and techniques failed to address structural issues in D3FEND’s core modeling approach, that were raised back in 2023[25], such as missing concepts, semantic overload, and a systematic lack of constraints that make the model under-specified.
+Although the new D3FEND extension represents significant practical improvements, OT users wan a primary representation for CTI exchange still need STIX, which D3FEND isn’t compatible with.
 Such contributions underscore the necessity of domain-specific CTI methodologies that address the unique operational and security challenges inherent in different CPSs.
@@
 Jin et al.
-[jin2024sharing] provided one of the first comprehensive empirical measurements of CTI sharing effectiveness, analyzing STIX volume, timeliness, coverage, and quality across public sources.
+[17] provided one of the first comprehensive empirical measurements of CTI sharing effectiveness, analyzing STIX volume, timeliness, coverage, and quality across public sources.
 Their findings reveal significant variations in feed quality and highlight the challenges that organizations encounter while operating in a shared intelligence environment.
-This empirical foundation complements broader systematic reviews that examine how organizations actually use CTI for security decision-making [ainslie2023cyber] and identify persistent barriers to effective CTI sharing, including governance issues, misaligned incentives, and insufficient standardization [alaeifar2024current].
+This empirical foundation complements broader systematic reviews that examine how organizations actually use CTI for security decision-making [2] and identify persistent barriers to effective CTI sharing, including governance issues, misaligned incentives, and insufficient standardization [3].
 Quality and trust of shared feeds remain persistent issues across the CTI ecosystem.
 Griffioen et al.
-[griffioen2019quality] conducted systematic quality evaluations of cyber threat intelligence feeds, revealing inconsistent timeliness and specificity that can undermine the operational value of shared intelligence.
+[16] conducted systematic quality evaluations of cyber threat intelligence feeds, revealing inconsistent timeliness and specificity that can undermine the operational value of shared intelligence.
 Similarly, Schaberreiter et al.
-[schaberreiter2019quantitative] developed quantitative frameworks for evaluating trust in CTI quality, demonstrating that trust and quality metrics vary significantly across different sources and feed types.
+[31] developed quantitative frameworks for evaluating trust in CTI quality, demonstrating that trust and quality metrics vary significantly across different sources and feed types.
 Together, these studies highlight the need for more rigorous quality assurance mechanisms in CTI sharing initiatives.
@@
 Standardization efforts play a crucial role in making threat observables portable and actionable across different organizational contexts.
-STIX 2.1, approved as an OASIS Standard (on 10 June 2021), serves as the exchange format used across many CTI producers and consumers [stix_v21_official].
+STIX 2.1, approved as an OASIS Standard (on 10 June 2021), serves as the exchange format used across many CTI producers and consumers [24].
 However, challenges remain in effectively representing complex ICS-specific artifacts within existing STIX objects.
-Zych and Mavroeidis [zych2022enhancing] have proposed enhancements to MITRE ATT&CK representations by STIX, to improve filtering and prioritization tasks, while RFC 9424 [rfc9424] provides important clarification on the capabilities and limitations of IOCs, reinforcing the value of higher-fidelity observables for creating durable detection capabilities.
+Zych and Mavroeidis [39] have proposed enhancements to MITRE ATT&CK representations by STIX, to improve filtering and prioritization tasks, while RFC 9424 [26] provides important clarification on the capabilities and limitations of IOCs, reinforcing the value of higher-fidelity observables for creating durable detection capabilities.
 Recent advances in natural language processing have introduced new approaches for extracting threat intelligence from unstructured reports.
-While some pipelines pursue full STIX coverage through classical NLP techniques combined with external knowledge bases [marchiori2023stixnet], alternative approaches rely on large language models to extract observables directly from text without requiring extensive knowledge graphs.
+While some pipelines pursue full STIX coverage through classical NLP techniques combined with external knowledge bases [20], alternative approaches rely on large language models to extract observables directly from text without requiring extensive knowledge graphs.
 These automated extraction techniques show promise for scaling CTI processing, though systematic evaluation reveals significant limitations and failure modes.
 Alam et al.
-[alam2024ctibench] developed CTIBench, a comprehensive benchmark specifically designed to evaluate LLM performance in cyber threat intelligence tasks across multiple cognitive dimensions, including memorization, understanding, problem-solving, and reasoning.
+[4] developed CTIBench, a comprehensive benchmark specifically designed to evaluate LLM performance in cyber threat intelligence tasks across multiple cognitive dimensions, including memorization, understanding, problem-solving, and reasoning.
 Their evaluation of five state-of-the-art models revealed concerning patterns: all models struggled significantly with tasks requiring domain expertise, with even the best-performing model (GPT-4) achieving only 71% accuracy on foundational CTI knowledge questions.
@@
 Fieblinger et al.
-[fieblinger2024actionable] developed a comprehensive methodology for automating the extraction of actionable CTI using LLMs for generating Knowledge Graphs.
+[15] developed a comprehensive methodology for automating the extraction of actionable CTI using LLMs for generating Knowledge Graphs.
 Their approach systematically evaluates multiple open-source LLMs (including Llama 2 series, Mistral 7B Instruct, and Zephyr) across different extraction techniques, such as few-shot prompt engineering, guidance frameworks, and fine-tuning to optimize the extraction of meaningful triples from CTI texts.
@@
 Bhusal et al.
-[bhusal2024secure] developed the SECURE benchmark using MITRE ATT&CK as a foundation for generating cybersecurity LLM performance evaluation across extraction, understanding, and reasoning tasks.
+[7] developed the SECURE benchmark using MITRE ATT&CK as a foundation for generating cybersecurity LLM performance evaluation across extraction, understanding, and reasoning tasks.
 Their approach demonstrates the potential of leveraging structured threat intelligence frameworks like ATT&CK for creating systematic evaluation datasets.
@@
 These results underscore the importance of meticulous validation and human oversight when deploying LLMs for threat intelligence processing, particularly in critical infrastructure environments where inaccurate information can have severe consequences.
-2.2.1 ATT&CK Framework Application in OT Environments
-For technique-level grounding in OT environments, the MITRE ATT&CK framework and its ICS extension provide structured approaches to understanding adversary behaviors [mitre_attack_design, mitre_attack_ics_design].
+ATT&CK Framework Application in OT Environments.
+For technique-level grounding in OT environments, the MITRE ATT&CK framework and its ICS extension provide structured approaches to understanding adversary behaviors [35, 5].
 Afenu et al.
-[afenu2024industrial] have leveraged ATT&CK to validate ICS defense mechanisms and testbed implementations.
+[1] have leveraged ATT&CK to validate ICS defense mechanisms and testbed implementations.
 Choi et al.
-[choi2021probabilistic] generated realistic OT attack sequences based on analysis of real-world datasets.
+[9] generated realistic OT attack sequences based on analysis of real-world datasets.
 These applications demonstrate the framework’s utility in bridging the gap between theoretical threat models and practical security implementations.
+D3FEND Framework Extension for OT Environments.
+TBD
 The existing literature highlights several recurring challenges in CTI, including limited interoperability of detection tools, insufficient standardization of threat-sharing protocols, and a lack of actionable intelligence for specific adversarial behaviors.
 While Krasznay et al.
-[krasznay2021possibilities] emphasized the importance of legislative and collaborative frameworks, López-Morales et al.
-[lopez2024securing] focused on advancing technical methodologies to enhance the granularity and applicability of CTI.
+[18] emphasized the importance of legislative and collaborative frameworks, López-Morales et al.
+[19] focused on advancing technical methodologies to enhance the granularity and applicability of CTI.
 Both approaches converge on the need for tailored threat intelligence that integrates sector-specific requirements into practical and effective security solutions.
@@
 Solid arrows represent common existing information flows.
-Typically, governments facilitate information sharing by providing reports with threat or vulnerability information to critical infrastructures and other organizations [gao_cybersecurity_2023].
+Typically, governments facilitate information sharing by providing reports with threat or vulnerability information to critical infrastructures and other organizations [37].
 The National Cyber Emergency Response Teams (CERT) is an important information-sharing hub, as it has unique access to threat information through the facilitation of information-sharing programs with both federal, defense, and intelligence agencies, as well as partnerships with private sector organizations.
-Security vendors provide tools/platforms and information to support the detection and mitigation of cyber threats [splunk_tdir, crowdstrike_tdir].
+Security vendors provide tools/platforms and information to support the detection and mitigation of cyber threats [34, 10].
 Many of these vendors search for vulnerabilities and threat actor activity to develop analytics or rules to be incorporated into detection tools.
@@
 Further, many vendors obtain data from their detection infrastructures deployed at asset owner sites or collect information while supporting incident response efforts.
-ICS vendors typically support information sharing efforts through the disclosure of security advisories and patches when a vulnerability is discovered in their software or devices [ics_patch_tuesday].
+ICS vendors typically support information sharing efforts through the disclosure of security advisories and patches when a vulnerability is discovered in their software or devices [33].
 Most importantly, ICS vendors have unique knowledge regarding the technical details for their products, which are typically based on proprietary technology.
 Asset owners are responsible for managing the risk presented by the cyber threats.
-The larger organizations that maintain critical infrastructures typically have dedicated security operations teams, the smaller commercial asset owners commonly rely on their national CERT [Cybersecurity_Guide_for_SMEs].
+The larger organizations that maintain critical infrastructures typically have dedicated security operations teams, the smaller commercial asset owners commonly rely on their national CERT [27].
 They may have a person partially employed in security operations to comply with regulations.
@@
 For example, STIX standard provides a JSON format to communicate threat behaviors.
-This includes definitions of SCOs [stix_v21_official] for specific types of artifacts that can be used to detect malicious activities or collect information about them.
+This includes definitions of SCOs [24] for specific types of artifacts that can be used to detect malicious activities or collect information about them.
 Other formats include OpenIOC [^1] https://www.scribd.com/document/653654203/An-Introduction-to-OpenIOC and MISP XML [^2] https://www.misp-project.org/.
@@
 4.1.1 Triton
-The Triton malware [CISA2017HatMan], also known as TRISIS or HatMan, targeted safety instrumented systems within an oil refinery by manipulating both an engineering workstation and the Triconex safety controllers used in the sulfur recovery units and burn management systems.
+The Triton malware [11], also known as TRISIS or HatMan, targeted safety instrumented systems within an oil refinery by manipulating both an engineering workstation and the Triconex safety controllers used in the sulfur recovery units and burn management systems.
 While the broader campaign during this attack included activities used to gain access to the facility (references as TEMP.Veles in ATT&CK), the defined Tactics, Techniques and Proceduress specific to the ICS environment primarily address the behaviors associated with the manipulation of Windows-based engineering workstations, safety controllers, and the associated network communications between them.
@@
 The adversarial techniques using this protocol include discovering target devices [T0846], detecting and changing their operating mode [T0858] [T0868], downloading programs to the targeted safety controller [T0843], executing the program [T0871], and extracting programs from the device [T0845].
-Developing detections for this activity requires the ability to (i) understand and parse the TriStation protocol  [tristation_protocol_reversing], (ii) identify specific functions of the protocol, and (iii) represent function payloads (e.g., program downloads contents).
+Developing detections for this activity requires the ability to (i) understand and parse the TriStation protocol  [21], (ii) identify specific functions of the protocol, and (iii) represent function payloads (e.g., program downloads contents).
 The STIX Network Traffic observable could be used to represent this activity, specifically the src/dst_payload_ref objects, which reference specific strings within the protocol payloads.
@@
 This could potentially be detected by the Process: Hash STIX object, while many devices produce CRC’s or checksums of deployed programs, these could be easily converted to a hash to support this object type.
-These Cyclic Redundancy Checks within the Triconex devices are a proprietary but documented function [triconex_safety_evaluation].
+These Cyclic Redundancy Checks within the Triconex devices are a proprietary but documented function [38].
 Most of these other Triconex behaviors are difficult to detect because they manipulate the device’s run-time memory, which depends on the undocumented and proprietary operation of the device.
@@
 Since these vulnerabilities have been known to be exploited by threat actors, they present a clear risk to asset owner environments, requiring some mitigating actions.
-While within IT environments these vulnerabilities are typically patched, it is well documented asset owners struggle to quickly deploy patches [sans_ics_ot_cybersecurity].
+While within IT environments these vulnerabilities are typically patched, it is well documented asset owners struggle to quickly deploy patches [28].
 If an asset owner cannot quickly patch these vulnerabilities to mitigate their risk, they may need to deploy detection mechanisms to prevent their exploitation.
@@
 Therefore, mechanisms to support understanding program segments on the controller, both through source and compiled code, may be effective in detecting malicious logic.
-    • Integrity checks: Further, many controllers also use integrity check mechanisms, typically based on CRCs or Checksums, to verify that programs have not changed  [plc_security_practices].
+    • Integrity checks: Further, many controllers also use integrity check mechanisms, typically based on CRCs or Checksums, to verify that programs have not changed  [30].
 On the one hand, security products can verify that a specific program has been downloaded to the device by validating its signatures.
-On the other hand, CRC/Checksum integrity checks are not sufficient against threat actors who intentionally modify the PLC code  [schuett2014evaluation].
-There is a need for a wider deployment of secured integrity signatures  [banerjee2024designing].
+On the other hand, CRC/Checksum integrity checks are not sufficient against threat actors who intentionally modify the PLC code  [32].
+There is a need for a wider deployment of secured integrity signatures  [6].
     • ICS Protocols: Currently, the STIX standard supports prevalent IT protocols, including HTTP and SMTP.